# An effective $\boldsymbol{\Lambda}$-Szekeres modelling of the local Universe with Cosmicflows-4

### v1

@@ Appendix B Code routine validation
$\mathop{\text{d}\!}s^{2}=-\mathop{\text{d}\!}t^{2}+\frac{R^{\prime 2}(t,r)}{1-k(r)}\mathop{\text{d}\!}r^{2}+R^{2}(t,r)\mathrm{d}^{2}\Omega\;,$ (B.2)
we can directly employ the redshift as parameter along null geodesics, and obtain the radial geodesic equations [Krasiński_1997, Celerier_2024]
$\frac{dt}{dz}=-\frac{R^{\prime}(r,t)}{(1+z)\dot{R}^{\prime}(r,t)}\;,\;\;\;\frac{dr}{dz}=-\frac{\sqrt{1-k(r)}}{(1+z)\dot{R}^{\prime}(r,t)}\;,$ (B.3)
with the luminosity and angular distances then given by [Krasiński_1997, Celerier_2024]
$d_{A}=\frac{d_{L}}{1+z^{2}}=R(r,t)\;.$ (B.4)
@@
$\chi(t^{*},r):=\int_{0}^{r}\frac{R^{\prime}(t^{*},r)}{\sqrt{1-k(\tilde{r})}}\mathop{\text{d}\!}\tilde{r}\;$. (B.5)
The latter can also be related to the l.o.s. comoving distance corresponding to an observed redshift $z$ via
$\chi(z)=\int_{0}^{z}\frac{c\mathop{\text{d}\!}z^{\prime}}{(1+z^{\prime})({\dot{R}^{\prime}}/{R^{\prime}})}\;.$ (B.6)
For the Hubble Bubble model, Eq. (B.6) becomes
$\chi(z)=c\int_{0}^{\rm{min}(z,z_{b})}\frac{dz^{\prime}}{\bar{H}+\delta H}+u(z-z_{b})c\int_{z_{b}}^{z}\frac{dz^{\prime}}{\bar{H}}$ (B.7)
@@
$r_{b}=c\int_{0}^{z_{b}}\frac{\mathop{\text{d}\!}z}{\bar{H}+\delta H}\;.$ (B.8)
Notice that in the simulation we consider distances in the comoving grid (Mpc/h) with respect to the outer flat FLRW region, hence $h=H/100$. However, the “true” radius should be given by $\tilde{h}=(\bar{H}+\delta H)/100$, so that
$r_{b}=\frac{Hr_{grid}}{H+\delta H}\;,$ (B.9)
@@
$z_{b}=(\bar{H}+\delta H)\frac{r_{b}}{c}\;,$ (B.10)
that together with Eq. (B.7) can be used to define our “truth” for the Hubble diagram. To validate the code routine, we thus test the comoving distance predictions of the various effective models with respect to the “truth” of the underlying Hubble bubble model.
Here, we consider an FLRW cosmological background with $H_{0}=74~\mathrm{km\,s^{-1}\,Mpc^{-1}}$ and $\Omega_{M0}=0.3$, consistent with the late-time results from CF4 analyses [Duangchan_2025]. We then take a central overdensity with linear density contrast of $\delta=0.1$, a radius within the comoving grid of $r_{b}\approx 80~\mathrm{Mpc/h}$, and a variation in the internal expansion rate of $\delta H\approx-1.2~\mathrm{km\,s^{-1}\,Mpc^{-1}}$. Then, from the density contrast, we directly compute the PVs and the Newtonian potential within the grid by applying linear perturbation theory, whilst imposing FLRW junction conditions at the grid boundaries (see the discussion in subSec. 2.1, and the reviews [KodamaSasaki_1984, Mukhanov_1992, Malik_2009]). In Fig. 7 we show on the SGX-SGY plane the density contrast, the variation in the internal Hubble parameter, the normalised Newtonian potential ($\Phi_{\mathrm{N}}/c^{2}$), and the radial PVs within our simulation. As expected, we see that the predicted radial PVs—induced by the Newtonian gravitational potential—extend beyond the spherical overdensity.
@@
We follow the procedure detailed in Sec. 2 to fit a $\Lambda$-Szekeres model to the simulated spherical overdensity and extract the quasilocal expansion field. We then define an approximate comoving distance for the effective $\Lambda$-Szekeres model—in accordance with the discussion in Sec. 3—expressed in redshift space as
$\chi_{\Lambda-\mathrm{Sk}}(z):=c\int_{0}^{z}\frac{\mathop{\text{d}\!}z}{H_{q}(z)}\,.$ (B.11)
@@
$\chi_{\mathrm{FLRW}}(z):=\frac{cz}{\bar{H}}\,$, (B.12)
@@
$\chi_{\mathrm{NPT}}(z):=\frac{cz}{\bar{H}}-\frac{v_{r}(z)}{\bar{H}}-\frac{c\Delta\Phi_{\mathrm{N}}(z)}{\bar{H}}\,$, (B.13)
@@
Figure 8: Percent differences, relative to the underlying Hubble-model comoving distance, of the predicted comoving distances for the $\Lambda$-Szekeres model, the cosmological FLRW background, and first-order perturbation theory around this background.
To conclude, we have simulated the Hubble Bubble model—a fully analytical, nonlinear effective model for overdensities over an FLRW background—and compared the comoving distance predictions for luminous sources obtained with our $\Lambda$-Szekeres reconstruction code to those of the model itself, as well as to the pure FLRW background and linear perturbation theory around it. We find that the predictions from our code closely match the true distances, outperforming linear perturbation theory within the Bubble while remaining comparable to it outside. Hence, the $\Lambda$-Szekeres reconstruction ultimately provides a fairly accurate method for modelling distance measures in inhomogeneous cosmological settings.

### v2

@@ Appendix C Code routine validation
$\mathop{\text{d}\!}s^{2}=-\mathop{\text{d}\!}t^{2}+\frac{R^{\prime 2}(t,r)}{1-k(r)}\mathop{\text{d}\!}r^{2}+R^{2}(t,r)\mathrm{d}^{2}\Omega\;,$ (C.2)
we can directly employ the redshift as a parameter along null geodesics, and obtain the radial geodesic equations [Krasiński_1997, 19]
$\frac{dt}{dz}=-\frac{R^{\prime}(r,t)}{(1+z)\dot{R}^{\prime}(r,t)}\;,\;\;\;\frac{dr}{dz}=-\frac{\sqrt{1-k(r)}}{(1+z)\dot{R}^{\prime}(r,t)}\;,$ (C.3)
with the luminosity and angular distances then given by [Krasiński_1997, 19]
$d_{A}=\frac{d_{L}}{1+z^{2}}=R(r,t)\;.$ (C.4)
@@
$\chi(t^{*},r):=\int_{0}^{r}\frac{R^{\prime}(t^{*},r)}{\sqrt{1-k(\tilde{r})}}\mathop{\text{d}\!}\tilde{r}\;$. (C.5)
This can also be related to the l.o.s. comoving distance corresponding to an observed redshift $z$ via
$\chi(z)=\int_{0}^{z}\frac{c\mathop{\text{d}\!}z^{\prime}}{(1+z^{\prime})({\dot{R}^{\prime}}/{R^{\prime}})}\;.$ (C.6)
For the Hubble Bubble model, Eq. (C.6) becomes
$\chi(z)=c\int_{0}^{\rm{min}(z,z_{b})}\frac{dz^{\prime}}{\bar{H}+\delta H}+u(z-z_{b})c\int_{z_{b}}^{z}\frac{dz^{\prime}}{\bar{H}}$ (C.7)
@@
$r_{b}=c\int_{0}^{z_{b}}\frac{\mathop{\text{d}\!}z}{\bar{H}+\delta H}\;.$ (C.8)
Notice that in the simulation, we consider distances in the comoving grid (Mpc/h) with respect to the outer flat FLRW region, hence $h=H/100$. However, the “true” radius should be given by $\tilde{h}=(\bar{H}+\delta H)/100$, so that
$r_{b}=\frac{Hr_{grid}}{H+\delta H}\;,$ (C.9)
@@
$z_{b}=(\bar{H}+\delta H)\frac{r_{b}}{c}\;,$ (C.10)
that together with Eq. (C.7) can be used to define our “truth” for the Hubble diagram. To validate the code routine, we thus test the comoving distance predictions of the various effective models with respect to the “truth” of the underlying Hubble bubble model.
Here, we consider an FLRW cosmological background with $H_{0}=74~\mathrm{km\,s^{-1}\,Mpc^{-1}}$ and $\Omega_{M0}=0.3$, consistent with the late-time results from CF4 analyses [37]. We then take a central overdensity with linear density contrast of $\delta=0.1$, a radius within the comoving grid of $r_{b}\approx 80~\mathrm{Mpc/h}$, and a variation in the internal expansion rate of $\delta H\approx-1.2~\mathrm{km\,s^{-1}\,Mpc^{-1}}$. Then, from the density contrast, we directly compute the PVs and the Newtonian potential within the grid by applying linear perturbation theory, whilst imposing FLRW junction conditions at the grid boundaries (see the discussion in Subsec. 2.1, and the reviews [62, 80, 72]). In Fig. 12 we show on the SGX-SGY plane the density contrast, the variation in the internal Hubble parameter, the normalised Newtonian potential ($\Phi_{\mathrm{N}}/c^{2}$), and the radial PVs within our simulation. As expected, we see that the predicted radial PVs induced by the Newtonian gravitational potential extend beyond the spherical overdensity.
@@
We follow the procedure detailed in Sec. 2 to fit a $\Lambda$-Szekeres model to the simulated spherical overdensity and extract the quasilocal expansion field. We then define an approximate comoving distance for the effective $\Lambda$-Szekeres model, in accordance with the discussion in Sec. 3, expressed in redshift space as
$\chi_{\Lambda-\mathrm{Sk}}(z):=c\int_{0}^{z}\frac{\mathop{\text{d}\!}z}{H_{q}(z)}\,.$ (C.11)
@@
$\chi_{\mathrm{FLRW}}(z):=\frac{cz}{\bar{H}}\,$, (C.12)
@@
$\chi_{\mathrm{NPT}}(z):=\frac{cz}{\bar{H}}-\frac{v_{r}(z)}{\bar{H}}-\frac{c\Delta\Phi_{\mathrm{N}}(z)}{\bar{H}}\,$, (C.13)
@@
Figure 13: Percent differences, relative to the underlying Hubble-model comoving distance, of the predicted comoving distances for the $\Lambda$-Szekeres model $\chi_{\rm\scriptscriptstyle\Lambda-Sk}$, the cosmological FLRW background $\chi_{\rm\scriptscriptstyle FLRW}$, and first-order perturbation theory $\chi_{\rm\scriptscriptstyle NPT}$ around this background.
To conclude, we have simulated the Hubble bubble —– a fully analytical, nonlinear effective model for overdensities over an FLRW background —– and compared the comoving distance predictions for luminous sources obtained with our $\Lambda$-Szekeres reconstruction code to those of the model itself, as well as to the pure FLRW background and linear perturbation theory around it. We find that the predictions from our code closely match the true distances, outperforming linear perturbation theory within the bubble while remaining comparable to it outside. Hence, the $\Lambda$-Szekeres reconstruction ultimately provides a fairly accurate method for modelling distance measures in inhomogeneous cosmological settings.